OpenStudy (anonymous):

Find b and c so that y = -20x^2+bx+c has vertex (-10, 3)

hero (hero):

The vertex form of a parabola is \(y = a(x - h)^2 + k\) We are given that (h,k) = (-10,3) and a = -20 so \(y = -20(x + 10)^2 + 3\) which expands to: \[y = -20(x + 10)(x + 10) + 3 \\= -20(x^2 + 20x + 100) + 3 \\ = -20x^2 - 400x - 2000 + 3 \\ = -20x^2 - 400x - 1997\] Therefore b = -400 and c = -1997 To check: http://www.wolframalpha.com/input/?i=-20x%5E2+-+400x+-1997+%3D+-20%28x+%2B+10%29%5E2+%2B3

OpenStudy (anonymous):

Thanks. That helped me greatly to understand the solution.

OpenStudy (anonymous):

I just checked and the answers that I got were b = -400 and c = -2003... my steps:

OpenStudy (anonymous):

-b / 2a = -b / 2(-20) = -b / -40 = - 10(-40) -b = 400 b = 400 -3 = -20(-10)^2+-400*-10+c -3 = -2000 + 4000 + c -3 = 2000 + c c = -2003

OpenStudy (anonymous):

@Hero

hero (hero):

If you have doubts check with wolframalpha